OpenStudy (anonymous):

how do you write the equation of the line passing through the points (-3,-12), (5,4)

OpenStudy (anonymous):

Do you know how to find the slope given 2 points?

OpenStudy (anonymous):

Oh and point-slope, or slope-int?

OpenStudy (anonymous):

y2-y1 over x2-x1?

OpenStudy (anonymous):

Yep thats right

OpenStudy (hba):

\[y-y_1=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\] Where (x_1,y_1)=(-3,-12) (x_2,y_2)=(5,4)

OpenStudy (anonymous):

just do that and ill get the y-mx+b? from that?

OpenStudy (anonymous):

No, hold on, ill write it step by step

OpenStudy (hba):

@cralls18 Just use that formula :)

OpenStudy (anonymous):